OpenStudy (anonymous):

what is the best way to integrate 1/sqrt(e^(2x)-25)?

OpenStudy (freckles):

\[\int\limits_{}^{}\frac{1}{ \sqrt{e^{2x}-25}} dx\] Hmmm... so we are looking at this...

OpenStudy (jhannybean):

double substitution?

OpenStudy (jhannybean):

u = 2x , du = 2dx \[\frac{1}{2}\int \frac{1}{\sqrt{e^u-25}}du\]

OpenStudy (freckles):

\[5 \sec(\theta)=e^x \text{ this might work too }\]

OpenStudy (jhannybean):

And then... p = e\(^u\) , dp = e\(^u\)du \(\implies\) du = \(\frac{dp}{e^u}\) \[\frac{1}{2}\int \frac{1}{\sqrt{p-25}}\cdot \frac{dp}{p}\]

OpenStudy (anonymous):

\[1/2\int\limits1/(\sqrt(p)(p+25)) dp\]

OpenStudy (anonymous):

i don't understand the step between your last response and this... I understand up to this point.

OpenStudy (anonymous):

Thanks for your help, btw.

OpenStudy (anonymous):

Freckles, how are you doing that substitution?

OpenStudy (freckles):

\[5 \sec(\theta)=e^x \\ 5 \sec(\theta) \tan(\theta)d \theta=e^x dx \text{ differentiated both sides } \\ \text{ since } e^x=5 \sec(\theta) \text{ we have } e^x \tan(\theta) d \theta=e^x dx \\ \text{ dividing both sides by } e^x \text{ gives } \\ \tan(\theta) d \theta=d x \\ \int\limits_{}^{} \frac{ \tan(\theta) d \theta}{\sqrt{(5 \sec(\theta))^2-25}}=\int\limits \frac{ \tan(\theta) d \theta}{\sqrt{25 \sec^2(\theta)-25}}\]

OpenStudy (freckles):

\[=\int\limits \frac{\tan(\theta)}{\sqrt{25} \sqrt{\sec^2(\theta)-1}} d \theta \] I think this can be finished from here pretty easily

OpenStudy (jhannybean):

My way is nearly the exact same way as freckles' but he just used trig subs to reduce the number of variable substitions in order to simplify the problem.

OpenStudy (jhannybean):

substitutions*